OpenStudy (anonymous):

Question 10 Given the functions f(x) = -2x - 1 and g(x) = -3x + 4, which operation results in the smallest coefficient on the x term? Two operations result in the same coefficient Subtraction Multiplication Addition Question 11 What is the simplified form of the fifth root of x times the fifth root of x times the fifth root of x times the fifth root of x? 4 times the fifth root of x the fifth root of 4x x to the five fourths power x to the four fifths power

OpenStudy (anonymous):

@math&ing001

OpenStudy (math&ing001):

You'll have to do it manually. I mean you try the subtraction, multiplication, and addition, then you compare the coefficient of the x term for each operation.

OpenStudy (math&ing001):

I can check you answer when you're done !

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

is it addition?

OpenStudy (math&ing001):

What did you get for multiplication ?

OpenStudy (anonymous):

i didnt do that one...

OpenStudy (math&ing001):

You should :)

OpenStudy (anonymous):

im confused.../;

OpenStudy (triciaal):

FOR QUESTION 11 when you multiply numbers with the same base add the powers base = x fifth root same as power = 1/5

OpenStudy (math&ing001):

f(x)*g(x) = 6x^2 - 5x - 4

OpenStudy (math&ing001):

Isn't it the same coefficient of x you got for addition ?

OpenStudy (math&ing001):

-5

OpenStudy (anonymous):

yes would that be right?

OpenStudy (math&ing001):

Yeah ! and you have it as an option

OpenStudy (anonymous):

so thats the final answer?

OpenStudy (math&ing001):

Yep

OpenStudy (anonymous):

YESSS!

OpenStudy (anonymous):

11?

OpenStudy (math&ing001):

x*x*x*x = x^4 and use what @triciaal said before.

OpenStudy (anonymous):

would it be the second one?

OpenStudy (math&ing001):

It's \[x ^{1/5}x ^{1/5}x ^{1/5}x ^{1/5} = x ^{\frac{ 4 }{ 5 }}\]

OpenStudy (triciaal):

no, you have x^(1/5) 4 times and you are multiplying the same base 4 times1/5 = 4/5 as shown above

OpenStudy (triciaal):

the last one